OpenStudy (anonymous):

during the first two semester 5 students withdrew from haileys algebra class.8 students were added to the class and 4 students were dropped as no shows by how many students did the original class size change

OpenStudy (anonymous):

say x is the total sice of the class... now x-5(dropped)+8(added)-4(dropped)= x-5+8-4=x-1 so one less then original size

OpenStudy (anonymous):

hey thnks